OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = 1/4; y intercept at (0, 1) slope = 1/4; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (anonymous):

Slope-Intercept form. y = mx + b m is slope b is y-intercept Change x - 4y = 4 into slope-intercept form~ x - 4y = 4 Subtract x from both sides. -4y = -x + 4 Divide both sides by -4. y = 1/4x - 1 So, what is your slope? What is the y-intercept?

OpenStudy (anonymous):

slope = 1/4; y intercept at (0, 1) is that correct?

OpenStudy (anonymous):

Very close: the y-intercept is -1

OpenStudy (anonymous):

Oh okay thats what I thought at first. But thanks a million.

OpenStudy (anonymous):

You're welcome. (: